kenzie126:

For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. A- 0.08 B- 0.496 C- 2.48 D- 5

AZ:

first plug in x = 3, what do you get? \( f(3) = 0.01(2)^3\)

AZ:

first do the exponents, what is 2^3 = and then multiply by 0.01

kenzie126:

idk

AZ:

\( 2^3\) is the same thing as 2 * 2 * 2 what is 2 * 2 * 2 = 2 times 2 times 2 =

kenzie126:

12

AZ:

what is 2 times 2

kenzie126:

i mean its 6. srry i was going by 4s lmaoo

AZ:

are you sure? what is 2 times 2

kenzie126:

4

AZ:

okay good, and what is 4 times 2

kenzie126:

8

AZ:

good now multiply that by 0.01

AZ:

you can use a calculator

AZ:

we then similarly have to do the same thing by plugging in x = 8 \(f(8) = 0.01(2)^8\)

AZ:

and we need to then find the average rate of change or the slope I'll throw in colors for you \(\Large\text{Slope} = \dfrac{\color{green}{y_2} - \color{orange}{y_1}}{\color{cyan}{x_2}-\color{red}{x_1}}\) where you have two points \(\Large (\color{red}{x_1}, \color{orange}{y_1})\) and \(\Large (\color{cyan}{x_2}, \color{green}{y_2})\)

AZ:

one of our points is (3, 0.08) what value do you get when x = 8?